Demand contracts from 500 to 400 when price rises from $40 to $42


1
Expert's answer
2022-03-17T09:39:01-0400

"PED= \frac{\Delta Q}{\Delta P}×\frac {P}{Q}"

"= \frac{(400-500)}{(42-40)}×\frac {P}{Q}"

"= \frac{-100}{2}×\frac {40}{500}= -4"